What is the minimum energy required to launch a satellite of mass m from the surface of a planet of mass M and radius R in a circular orbit at an altitude of 2R

Grade:12

1 Answers

Avishek Bhagat
3 years ago
Dear Student,
the answer is given as


E = (Energy of satellite) - (the energy of mass on the surface of a planet)

=−GMm/2r−(−GMm/R)=−GMm/2r−(−GMm/R)

Here, r = R + 2R = 3R

Substituting in about equation we get.
E= -GMm/6R+GMm/R
E=(5GMm)/6R

askIITians Expert
Yuvraj Singh